OpenStudy (anonymous):

x^1/2 +sqrt3x^1/4 -18 then, (x^1/4)^2 +sqrt3x^1/4 -18 then you factor, ((x^1/4)+ or - ?) repeat. How do you factor 18 to also when equal the sqrt3x

OpenStudy (dumbcow):

it doesn't...with an irrational coefficient you have to use quadratic formula

OpenStudy (psymon):

I'd copy and paste the quadratic formula of this that I posted on your previous question, but not sure how O.o

OpenStudy (dumbcow):

you can...highlight the latex, then right click, select maText i think .... then it opens in text format so you can copy it

OpenStudy (psymon):

\[\frac{ -(3^{\frac{ 1 }{ 4 }})\pm \sqrt{(3^{\frac{ 1 }{ 4 }})^{2}-4(-18)} }{ 2 }\]

OpenStudy (psymon):

Oh, cool :D

OpenStudy (anonymous):

so for inside the sqrt i got 30.9375, what next

OpenStudy (psymon):

Not sure how you got that inside of the square root O.o But since all of this is calculator work really anyway. When you do the whole thing, you get the answers 3.64 for the + part of it and -4.95 for the - part of it. So then you have this: \[x ^{\frac{ 1 }{ 4 }} = 3.64\] \[x ^{\frac{ 1 }{ 4 }}= -4.95\] Now ya just solve for x

OpenStudy (anonymous):

solve for x how?

OpenStudy (psymon):

Since x is to the 1/4 power, you would need to raise your answers to the 4th power.

OpenStudy (psymon):

Now what I should definitely point out is that when you take an even root of something, you cannot end up with a negative answer. So there is no possible way for x^(1/4) to = -4.95 or negative anything. So your one answer will be (3.64)^4

OpenStudy (anonymous):

\[x ^\frac{ 1 }{ 2 } -5\sqrt{3}x ^\frac{ 1 }{ 4 } +18\]

OpenStudy (anonymous):

they changed the problem on me for taking too long

OpenStudy (psymon):

Ah. Well, we had the last problem done anyway, so this problem is the same thing anyway :/ gimme a sec for calculator magic, haha x_x

OpenStudy (anonymous):

great ! cause i dont have one

OpenStudy (psymon):

So wait, is the 3 also to the 1/4 power or only square root?

OpenStudy (anonymous):

just how its written

OpenStudy (psymon):

Okay, so square root 3 then it looks like and only the x is to the 1/4 power.

OpenStudy (anonymous):

i believe its to the sqrt

OpenStudy (psymon):

144 and 729.

OpenStudy (anonymous):

3 is sqrt, x and 1/4 are not in the sqrt

OpenStudy (psymon):

Yeah, thats how I solved it then, so those are the two answers.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Want to help me with a few others since you have a calculator? (:

OpenStudy (anonymous):

It would me a lot to me(:

OpenStudy (anonymous):

mean*

OpenStudy (psymon):

Well, if they're ones I can go through an explaining process with then I'd like to do that, but if it's just ridiculous things like that then yeah, doesnt bug me, lol.

OpenStudy (psymon):

Geez, finally working again

OpenStudy (anonymous):

i know. \[x+\sqrt{x}-2\]

OpenStudy (anonymous):

im running out of time . lol

OpenStudy (psymon):

solve for x again, right?

OpenStudy (anonymous):

yeah ,quadratic form

OpenStudy (psymon):

Now this one we CAN do by normal factoring. Let sqrtx = u. Which then gives us: \[u ^{2}+u -2 = 0\]

OpenStudy (psymon):

Easy one to factor here :P (u-1)(u+2)

OpenStudy (anonymous):

(3x)^2-12(3x)+27

OpenStudy (psymon):

Oh, -9 and -3, that one is easy, too o.o

OpenStudy (psymon):

I mean (x-9)(x-3) where x = 3x xD

OpenStudy (anonymous):

oh well i just put the -9,-3

OpenStudy (psymon):

Oh x_x

OpenStudy (anonymous):

\[x-2\sqrt{x}-35\]

OpenStudy (anonymous):

no solution, 49, 25,49, 25i,49 these are the choices